# GenMix: Combining Generative and Mixture Data Augmentation for Medical Image Classification

### v1

@@ 1 Introduction
Data augmentation (DA) plays a crucial role in medical image analysis, significantly enhancing the performance of machine learning models by increasing the diversity and volume of training data (Shorten & Khoshgoftaar, 2019). This is particularly important in medical imaging, where acquiring large and varied datasets is often challenging due to privacy concerns, high costs, and the rarity of certain conditions. Effective DA techniques can help mitigate issues such as over-fitting and class imbalance, thereby improving model generalization and robustness. However, standard augmentation methods such as rotation, scaling, and flipping may not be sufficient to capture the complex variations and subtle differences present in medical images, necessitating the development of more sophisticated approaches.
@@
In recent years, generative models like Generative Adversarial Networks (GANs) and Diffusion Models have gained popularity in medical image analysis for their ability to create realistic synthetic data. These models can generate new patterns that resemble real medical images, thus expanding the training dataset. For instance, Salehinejad et al. used Deep Convolutional GAN (DCGAN) (Radford et al., 2015) and AlexNet (Krizhevsky et al., 2012) to improve classification performance in the task of identifying five diseases of chest X-ray images (Salehinejad et al., 2018). Frid-Adar et al. employed DCGAN and Auxiliary Conditional GAN (ACGAN) (Odena et al., 2016) to classify liver lesions in abdominal CT images using LeNet-like network (Frid-Adar et al., 2018). Zhao et al. proposed Forward and Background GAN (F&BGAN) to classify lung nodules in chest CT images using modified VGG-16 network (Zhao et al., 2018). Lee et al. used DCGAN and pix2pix (Isola et al., 2016) to classify focal liver lesions in abdominal CT images using AlexNet (Lee et al., 2021). Generative models offer the advantage of generating high-quality synthetic data of novel patterns, which can be more diverse and representative of real-world variability compared to transform-based data augmentation techniques. However, generative models often face challenges such as mode collapse in GANs and the complexity of training diffusion models, especially when the amount of available training data is limited.
@@
Alternatively, mixture models, which combine data from different classes, have been used to enhance class boundary regions, making models more robust to variations. For instance, Nishio et al. used MixUp (Zhang et al., 2017) to classify COVID-19, pneumonia, and normal chest X-ray images using the VGG-16 network (Nishio et al., 2020). Rahan et al. applied MixUp to classify five diseases in chest X-ray images using the ResNet-18 network (Rajan et al., 2021). Özdemir et al. employed MixUp to distinguish COVID-19 from normal chest CT images using the ResNet-50 and ResNet-101 networks (Özdemir & Sönmez, 2022). Mixture models have been shown to improve the generalization and robustness of deep learning models by diversifying the patterns of training data despite their simple calculations. However, mixture models can become biased towards major classes in scenarios of class imbalance, leading to suboptimal performance.
@@
To address the limitations of existing DA techniques, we propose a novel approach called GenMix. GenMix synergistically combines the strengths of generative and mixture models, leveraging the benefits of both methodologies while mitigating their individual shortcomings. Our approach consists of two main stages: (1) training a generative model to produce synthetic images, and (2) performing mixup between synthetic and real data. By integrating these steps, GenMix improves the quality and diversity of synthetic data, while simultaneously enhancing the learning of new data patterns from generative models and the boundary strengthening properties of mixture models. The experimental results demonstrate that the proposed GenMix consistently improves the effectiveness of data augmentation across different types of generative models. Moreover, a type of diffusion model known as Textual Inversion (Gal et al., 2023) achieved the best classification performance on the FLL dataset without any fine-tuning, solely by applying GenMix.
@@ 2.1 Synthetic Image Generation with GANs and Diffusion Models
In our proposed method, we use a generative model to generate synthetic images, specifically using DCGAN (Radford et al., 2015), StyleGAN (Karras et al., 2018), Diffusion-based Textual Inversion (Gal et al., 2023), and Diffusion Model (Dhariwal & Nichol, 2021), which are state-of-the-art models capable of generating synthetic images.
@@
Textual Inversion. The Textual Inversion is designed to utilize text-to-image generation diffusion models without retraining them on new datasets. It introduces ”textual tokens,” which serve as placeholders for specific subjects or objects within the model’s pre-existing vocabulary. During the optimization process, these tokens are fine-tuned to represent unique entities, allowing the model to generate images of these entities when provided with text prompt. This method leverages the robustness and generalizability of large-scale text-to-image models while providing a flexible and efficient means of generating customized images, tailored to represent specific subjects or styles not directly covered in the model’s original training data.
In the proposed method, we employ Textual Inversion to make Latent Diffusion Model (LDM) (Rombach et al., 2022) pre-trained on LAION-400B (Schuhmann et al., 2021) to generate FLL images without additional training. The Textual Inversion optimizes unique textual tokens serving as placeholders, which, through optimization, are adjusted to represent the specific characteristics of FLLs within the pre-trained framework of the diffusion model. Consequently, when these optimized tokens are included in a text prompt, the model generates synthetic images that mimic the desired features of FLLs, thereby enabling a customized data augmentation strategy without the computational overhead and data requirements typically associated with model finetuning. The primary advantage of generating synthetic FFL images using Textual Inversion lies in its efficiency and flexibility. By leveraging the capabilities of a pre-trained diffusion model, it provides a practical solution for augmenting FFL datasets without the need for extensive computational resources and large datasets. However, the quality and diversity of the generated images depend on the capabilities and biases of the pre-trained model. If the model has not encountered sufficient variations of FLL-like images during its initial training, the synthetic images may not fully capture the range of appearances that real FLLs can exhibit.
@@
In the proposed method, we trained ImageNet-pretrained ADM diffusion model (Dhariwal & Nichol, 2021) on a training FLL dataset, and generated the synthetic images. Fig. 2 shows the evolution of the synthetic FLL images generative by ADM during the training process. Initially, the generated images were blurry and lacked distinct features, typical of early training phases in diffusion models. As training progressed, the images began to exhibit clearer structures and more defined features, indicating that the model was learning the underlying patterns of FLL images. By the final stages of training, the synthetic images were nearly indistinguishable from real FLL images.
@@ 3.1 Experimental Settings
Datasets. We conducted an experiment on the task of classifying diseases of FLLs on abdominal CT images (Bae et al., 2021; Lee et al., 2021). The dataset consists of CT scans collected from 502 colorectal cancer patients, including 1,290 focal liver lesion patches with 676 cysts, 130 hemangiomas, and 484 metastases. All images have a resolution of $512\times 512$, a pixel size in the range of $0.5\times 0.5$ mm2 to $0.8\times 0.8$ mm2, and slice thickness between 3 to 5 mm. All FLLs are manually segmented on the axial plane with the largest cross section and the dataset is split into a training set of 681 lesions (433C+70H+178M), a validation set of 302 lesions (115C+30H+157M), and a test set of 307 lesions (128C+30H+149M) based on the acquisition date.
@@
Implementation Details. For training DCGAN and StyleGAN, we set the hyperparameters to 70,000 iterations, a batch size of 8, and a learning rate of 0.001. The number of generated images was set to be equal to the number of training images for each class. In synthetic image generation stage, we employed the LDMs (Rombach et al., 2022) pre-trained on the LAION-400M dataset (Schuhmann et al., 2021) for applying Textual Inversion. In MixUp stage of the proposed method, we set the MixUp parameter to $\alpha=0.3$. We used the ImageNet-pretrained VGG-16 (Simonyan & Zisserman, 2014) network for the FLL classification. For training VGG-16, we set the hyperparameters to 100 epochs, a batch size of 8, a learning rate of 0.0002, and an early stopping condition of 20.

### v2

@@ 1 Introduction
Data augmentation (DA) plays a crucial role in medical image analysis, significantly enhancing the performance of machine learning models by increasing the diversity and volume of training data [1]. This is particularly important in medical imaging, where acquiring large and varied datasets is often challenging due to privacy concerns, high costs, and the rarity of certain conditions. Effective DA techniques can help mitigate issues such as over-fitting and class imbalance, thereby improving model generalization and robustness. However, standard augmentation methods such as rotation, scaling, and flipping may not be sufficient to capture the complex variations and subtle differences present in medical images, necessitating the development of more sophisticated approaches.
@@
In recent years, generative models like Generative Adversarial Networks (GANs) and Diffusion Models have gained popularity in medical image analysis for their ability to create realistic synthetic data. These models can generate new patterns that resemble real medical images, thus expanding the training dataset. For instance, Salehinejad et al. used Deep Convolutional GAN (DCGAN) [2] and AlexNet [3] to improve classification performance in the task of identifying five diseases of chest X-ray images [4]. Frid-Adar et al. employed DCGAN and Auxiliary Conditional GAN (ACGAN) [5] to classify liver lesions in abdominal CT images using LeNet-like network [6]. Zhao et al. proposed Forward and Background GAN (F&BGAN) to classify lung nodules in chest CT images using modified VGG-16 network [7]. Lee et al. used DCGAN and pix2pix [8] to classify focal liver lesions in abdominal CT images using AlexNet [9]. Generative models offer the advantage of generating high-quality synthetic data of novel patterns, which can be more diverse and representative of real-world variability compared to transform-based data augmentation techniques. However, generative models often face challenges such as mode collapse in GANs and the complexity of training diffusion models, especially when the amount of available training data is limited.
@@
Alternatively, mixture models, which combine data from different classes, have been used to enhance class boundary regions, making models more robust to variations. For instance, Nishio et al. used MixUp [10] to classify COVID-19, pneumonia, and normal chest X-ray images using the VGG-16 network [11]. Rahan et al. applied MixUp to classify five diseases in chest X-ray images using the ResNet-18 network [12]. Özdemir et al. employed MixUp to distinguish COVID-19 from normal chest CT images using the ResNet-50 and ResNet-101 networks [13]. Mixture models have been shown to improve the generalization and robustness of deep learning models by diversifying the patterns of training data despite their simple calculations. However, mixture models can become biased towards major classes in scenarios of class imbalance, leading to suboptimal performance.
@@
To address the limitations of existing DA techniques, we propose a novel approach called GenMix. GenMix synergistically combines the strengths of generative and mixture models, leveraging the benefits of both methodologies while mitigating their individual shortcomings. Our approach consists of two main stages: (1) training a generative model to produce synthetic images, and (2) performing mixup between synthetic and real data. By integrating these steps, GenMix improves the quality and diversity of synthetic data, while simultaneously enhancing the learning of new data patterns from generative models and the boundary strengthening properties of mixture models. The experimental results demonstrate that the proposed GenMix consistently improves the effectiveness of data augmentation across different types of generative models. Moreover, a type of diffusion model known as Textual Inversion [14] achieved the best classification performance on the FLL dataset without any fine-tuning, solely by applying GenMix.
@@ 2.1 Synthetic Image Generation with GANs and Diffusion Models
In our proposed method, we use a generative model to generate synthetic images, specifically using DCGAN [2], StyleGAN [15], Diffusion-based Textual Inversion [14], and Diffusion Model [16], which are state-of-the-art models capable of generating synthetic images.
@@
Textual Inversion. The Textual Inversion is designed to utilize text-to-image generation diffusion models without retraining them on new datasets. It introduces "textual tokens," which serve as placeholders for specific subjects or objects within the model’s pre-existing vocabulary. During the optimization process, these tokens are fine-tuned to represent unique entities, allowing the model to generate images of these entities when provided with text prompt. This method leverages the robustness and generalizability of large-scale text-to-image models while providing a flexible and efficient means of generating customized images, tailored to represent specific subjects or styles not directly covered in the model’s original training data.
In the proposed method, we employ Textual Inversion to make Latent Diffusion Model (LDM) [17] pre-trained on LAION-400B [18] to generate FLL images without additional training. The Textual Inversion optimizes unique textual tokens serving as placeholders, which, through optimization, are adjusted to represent the specific characteristics of FLLs within the pre-trained framework of the diffusion model. Consequently, when these optimized tokens are included in a text prompt, the model generates synthetic images that mimic the desired features of FLLs, thereby enabling a customized data augmentation strategy without the computational overhead and data requirements typically associated with model finetuning. The primary advantage of generating synthetic FFL images using Textual Inversion lies in its efficiency and flexibility. By leveraging the capabilities of a pre-trained diffusion model, it provides a practical solution for augmenting FFL datasets without the need for extensive computational resources and large datasets. However, the quality and diversity of the generated images depend on the capabilities and biases of the pre-trained model. If the model has not encountered sufficient variations of FLL-like images during its initial training, the synthetic images may not fully capture the range of appearances that real FLLs can exhibit.
@@
In the proposed method, we trained ImageNet-pretrained ADM diffusion model [16] on a training FLL dataset, and generated the synthetic images. Fig. 2 shows the evolution of the synthetic FLL images generative by ADM during the training process. Initially, the generated images were blurry and lacked distinct features, typical of early training phases in diffusion models. As training progressed, the images began to exhibit clearer structures and more defined features, indicating that the model was learning the underlying patterns of FLL images. By the final stages of training, the synthetic images were nearly indistinguishable from real FLL images.
@@ 3.1 Experimental Settings
Datasets. We conducted an experiment on the task of classifying diseases of FLLs on abdominal CT images [19, 9]. The dataset consists of CT scans collected from 502 colorectal cancer patients, including 1,290 focal liver lesion patches with 676 cysts, 130 hemangiomas, and 484 metastases. All images have a resolution of $512\times 512$, a pixel size in the range of $0.5\times 0.5$ mm2 to $0.8\times 0.8$ mm2, and slice thickness between 3 to 5 mm. All FLLs are manually segmented on the axial plane with the largest cross section and the dataset is split into a training set of 681 lesions (433C+70H+178M), a validation set of 302 lesions (115C+30H+157M), and a test set of 307 lesions (128C+30H+149M) based on the acquisition date.
@@
Implementation Details. For training DCGAN and StyleGAN, we set the hyperparameters to 70,000 iterations, a batch size of 8, and a learning rate of 0.001. The number of generated images was set to be equal to the number of training images for each class. In synthetic image generation stage, we employed the LDMs [17] pre-trained on the LAION-400M dataset [18] for applying Textual Inversion. In MixUp stage of the proposed method, we set the MixUp parameter to $\alpha=0.3$. We used the ImageNet-pretrained VGG-16 [20] network for the FLL classification. For training VGG-16, we set the hyperparameters to 100 epochs, a batch size of 8, a learning rate of 0.0002, and an early stopping condition of 20.